OpenStudy (anonymous):

Find the derivative. d/dx x^2 (integral) ln(7 + t^2) dt 0

OpenStudy (anonymous):

FTC along with chain rule

hartnn (hartnn):

\(\Large \dfrac{d}{dx}\int \limits_{f(x)}^{g(x)}h(x)dx = h(g(x))g'(x)-h(f(x))f'(x) \)

hartnn (hartnn):

\(\Large \dfrac{d}{dx}\int \limits_0^{g(x)}h(t)dt = h(g(t))g'(t) \) h(t) = ln(7+t^2)

hartnn (hartnn):

g(x) = x^2

OpenStudy (anonymous):

F(b)-F(a)… isn't that along the lines of what i could do? so i find the antiderivative and plug in those values right..?

OpenStudy (anonymous):

\[ \frac{d}{dx}\int\limits_{0}^{x^2}\ln(7+t^2)~dt = \frac{d}{dx}\int\limits_{0}^{x^2}g'(t)~dt = \frac{d}{dx}\Bigg[g(x^2)-g(0)\Bigg] \]

OpenStudy (anonymous):

To solve:\[ \frac{d}{dx}\Bigg[g(x^2)-g(0)\Bigg] \]You need to use the chain rule.

OpenStudy (anonymous):

so it'd be 1/(7+t^2) * 2t

OpenStudy (anonymous):

by the chain rule.. and then i plug in the values..?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

I didn't say \(g(t) = \ln(7+t^2)\). I said that \(g'(t) = \ln(7+t^2)\).

hartnn (hartnn):

\(\Large \dfrac{d}{dx}\int \limits_0^{g(x)}h(t)dt = h(g(x))g'(x)\) h(t) = ln (7+t^2) g(x) = x^2

OpenStudy (anonymous):

Also, the \(t\) went away when we integrated, so we won't be seeing it anymore.

OpenStudy (anonymous):

okay… so i need to find the antiderivative of g'(t)… I'm sorry my professor ddnt demonstrate how the chain rule was used for these problems

OpenStudy (anonymous):

i am so confused; im sorry

hartnn (hartnn):

the anti derivative of g'(t) is just g(t) which is shown by wio at this step \(\frac{d}{dx}\int\limits_{0}^{x^2}\ln(7+t^2)~dt = \frac{d}{dx}\int\limits_{0}^{x^2}g'(t)~dt = \frac{d}{dx}\Bigg[g(x^2)-g(0)\Bigg]\) in the end, he also plugged in the limits

hartnn (hartnn):

\(\frac{d}{dx}\int\limits_{0}^{x^2}\ln(7+t^2)~dt = \frac{d}{dx}\int\limits_{0}^{x^2}g'(t)~dt =\frac{d}{dx} [g(t)]_{0}^{x^2}= \frac{d}{dx}\Bigg[g(x^2)-g(0)\Bigg]\)

hartnn (hartnn):

and since g(0) is a constant, you just need \(\Large \dfrac{d}{dx}g(x^2)\) which will require chain rule

OpenStudy (anonymous):

The 'trick' is that \[ \frac{d}{dx}g(x^2) = \frac{d}{d(x^2)}g(x^2)\frac{d}{dx}x^2 = g'(x^2)(2x) \]

hartnn (hartnn):

and we call that trick as chain rule

OpenStudy (anonymous):

It is easier to see when you let \(u=x^2\), because then it is just: \[ \frac{dg}{dx}=\frac{dg}{du}\frac{du}{dx} \]

OpenStudy (anonymous):

I'm confused on why it requires the chain rule.. g(t) is composed with x^2 because of the whole F(b) -F(a), but i don't know why you just wouldn't stop at g(x^2). I understand how the chain rule works, I'm just not quite sure on why it takes place if that makes sense..?:/

OpenStudy (anonymous):

So these problems would be easier if i use the technique of integration by substitution

OpenStudy (anonymous):

The reason why we don't stop at \(g(t)\) is because we don't always know what \(g(t)\) is, but we know what \(g'(t)\) is, since it is the integrand.

OpenStudy (anonymous):

oooohhh okay! so what would be the following step?

OpenStudy (anonymous):

This was the furthest we got so far.\[ \frac{d}{dx}g(x^2) = \frac{d}{d(x^2)}g(x^2)\frac{d}{dx}x^2 = g'(x^2)\cdot 2x \]

OpenStudy (anonymous):

mhmm! i just don't know what to do with that. do we then apply the upper and lower limits?

OpenStudy (anonymous):

Well, we want to get rid of the \(g'\), by remember what it originally was.

OpenStudy (anonymous):

so g' was originally ln(7+t^2)

OpenStudy (anonymous):

Yep, and we need to sub in \(x^2\) for \(t\).

OpenStudy (anonymous):

so ln(7+(x^2)^2)

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

Do what is the whole answer?

OpenStudy (anonymous):

2x*ln(7+(x^2)^2)…?

OpenStudy (anonymous):

You can simplify \((x^2)^2\), can't you?

OpenStudy (anonymous):

thank you sooooo much!!!! i got it right :) they took the unsimplified version but it'd be x^4